# StoryBox: Collaborative Multi-Agent Simulation for Hybrid Bottom-Up Long-Form Story Generation Using Large Language Models

### v1

@@ Introduction
The advent of large language models (LLMs) has brought significant advancements to various fields, including multi-agent simulations (Li et al. 2024; Qian et al. 2024; Chen et al. 2024). These simulations offer a powerful tool for modeling complex interactions between virtual agents, providing a dynamic and context-rich environment for story generation. When humans write stories, they typically have an overarching mental picture of the story world. In our approach, multi-agent simulations are used to create this overarching scene, where agents (i.e., the characters in the story) interact based on predefined behaviors, triggering emergent events that form the foundation of the story. This allows for the automatic generation of diverse and engaging scenarios, which are crucial for building long-form stories.
@@
Traditional story generation often follows a top-down method (Zhou et al. 2023; Wang et al. 2023; Fan, Lewis, and Dauphin 2019; Goldfarb-Tarrant et al. 2020; Yang et al. 2022), where a story’s structure is outlined first and then expanded. Although this framework provides a guide, it can limit natural character development and plot progression, resulting in stories that feel forced or predictable. In contrast, our hybrid bottom-up approach starts with agent-driven simulations that generate emergent events. As agents interact, these events unfold naturally, enabling more organic character growth and plot evolution, ultimately producing stories with greater depth, coherence, and spontaneity.
@@
LLM-based multi-agent simulations have gained attention for their advanced language processing and decision-making, enabling nuanced agent interactions (Jinxin et al. 2023; Liu et al. 2023; Feng, Feng, and Qin 2023). These systems have been explored in various domains (Wang et al. 2024; Williams et al. 2023), particularly social simulation, which is most relevant to our research. Social simulation with LLM-based agents provides an effective way to model complex societal dynamics that are otherwise difficult or costly to study (Li, Yang, and Zhao 2023; Li et al. 2023). For example, S3 (Gao et al. 2023) investigates the spread of information, emotions, and attitudes in social networks, while Generative Agents (Park et al. 2023) and AgentSims (Lin et al. 2023) model daily human interactions in virtual towns. Social Simulacra (Park et al. 2022) focuses on simulating online community regulation, and SocialAI School (Kovač et al. 2023) uses LLMs to simulate child development. These studies highlight the growing potential of LLM-based agents to offer valuable insights into societal behavior, community regulation, and social development, which provides a foundation for hybrid bottom-up long-form story generation.
@@
Story generation has advanced with the rise of LLMs, which produce fluent stories but often struggle with coherence and consistency (Huot et al. 2024). To address these challenges, frameworks such as outlining followed by expansion into full stories have been proposed (Zhou et al. 2023; Wang et al. 2023). However, long-form story generation remains challenging. Recent studies have explored multi-agent systems using LLMs (Wang, Zhou, and Ledo 2024; Nasir, James, and Togelius 2024), such as Agents’ Room (Huot et al. 2024), which focuses on generating stories of 1,000-2,000 words using Planning and Writing Agents coordinated by an Orchestrator, and IBSEN (Han et al. 2024), which employs Director-Actor agent collaboration for scriptwriting. While top-down approaches, such as outlining, provide structure, they limit natural storylines and character interactions. In contrast, our hybrid bottom-up approach generates dynamic stories from simulations, producing richer and more coherent stories of over 10,000 words, with clear advantages over traditional methods.
@@
The environment plays a critical role in the sandbox. In Generative Agents, the environment is modeled using a tile-based system, where tiles represent the basic environmental units commonly found in RPG games. However, this approach has inherent limitations, particularly the need for precise coordinates, which can make modeling more rigid and less flexible. To overcome these limitations, we adopt a more general tree-like structure for environment modeling, which does not rely on specific coordinates but instead uses relative distances between objects and areas. For a detailed introduction to environment modeling, please refer to supplementary material.
@@
Following the setup described in DOC (Yang et al. 2023), we use premises, settings, and characters as inputs. However, the DOC dataset contains relatively homogeneous story types, so we constructed a new dataset consisting of 20 story-related settings without reference answers. Our dataset features a broader variety of story types, including genres such as science fiction. Detailed information about our dataset can be found in the supplementary material.
@@
We conduct structured pairwise comparisons across six criteria: Plot, Creativity, Character Development, Language Use, Conflict Quality, and Overall. These dimensions cover key aspects of long-form storytelling, including narrative coherence, originality, emotional depth, and stylistic quality. Annotators with backgrounds in literature, storytelling, and related fields compare story pairs from the same setting and provide ratings for each dimension as well as an overall preference. To reduce fatigue and bias, each annotator evaluates a limited number of pairs per session. We apply a Latin Square design to balance the assignment of story settings and randomize story order. For each setting, we collect $M\times N\times(N-1)/2$ comparisons, where $M$ is the number of settings and $N$ the number of methods.
@@
Full metric definitions, evaluation procedures, and interface details can be found in the supplementary material.
@@
We selected several high-performing open-source methods that can be reliably reproduced as our baselines. We categorize the baseline methods into three types: (1) Vanilla LLMs: methods that generate long-form stories directly using large language models, such as GPT-4o (Hurst et al. 2024) and DeepSeek-V3 (Liu et al. 2024); (2) Structured Frameworks: methods that utilize structured frameworks for long-form story generation based on LLMs, such as Re3 (Yang et al. 2022) and DOC-V2 (Yang et al. 2023); and (3) Multi-Agent Simulations: methods that generate through multi-agent simulations, such as IBSEN (Han et al. 2024).
@@ Conclusion
In this paper, we introduced StoryBox, a novel approach for long-form story generation using multi-agent simulations. Our experiments show that StoryBox outperforms existing methods on many metrics. Despite challenges in evaluating story quality, both automatic and human evaluations confirm its effectiveness in generating engaging and coherent stories. Future work will focus on refining the sandbox environment and exploring additional story-driven applications.
@@ Appendix C Metric Details
  • Character Behavior Consistency: A specialized metric for multi-agent simulation-based story generation, this measures whether characters’ actions align with their predefined persona settings (”Persona Scratch Information”). Characters should act consistently with their attributes, goals, and motivations, ensuring believability and coherence in the story.
@@ Appendix D Human Evaluation Details
To facilitate consistent and efficient story evaluation, we developed a custom web-based platform tailored to the needs of our human evaluation task. An overview of the platform interface is shown in Figure 8. To accommodate a diverse group of annotators, the platform supports both Chinese and English language modes. Access to the system is restricted to verified annotators, who must log in before participating in the evaluation process to ensure traceability and avoid duplicate submissions.
@@ Appendix F Sandbox Initialization
The setup of sandbox characters is derived from the understanding of the relevant story settings. It primarily involves establishing the “Persona Scratch Information”, as outlined in Figure 2, which includes details such as names, ages, and other relevant information. Based on the description in “Story 1”, we utilized GPT-4o mini to generate the corresponding character setups, as shown in Table LABEL:tab:story_1_character_setup. The same approach is applied to other stories.
@@
The sandbox environment setup is similarly based on the understanding of the relevant story settings. The goal is to construct an environment with hierarchical relationships, as depicted in Figure 7. In this case, drawing from the details in “Story 1”, we used GPT-4o mini to generate the corresponding environment setup in the form of a YAML file, as shown in Table LABEL:tab:story_1_environment_setup. This method is also applied to other stories in the dataset.
@@ Appendix H More Implementation Details
During the reproduction of DOC-V2, we observed that the prompts used in this method often resulted in parsing failures. Additionally, since this method does not specify the type of story, we set the story type of this method to “narrative” during evaluation. To ensure fairness, we replaced all OpenAI engines used in this method with GPT-4o mini.
@@
During the reproduction of IBSEN, we observed that this method primarily utilizes a multi-agent virtual sandbox for generating theatrical scripts. As a result, it is necessary to first convert the premise, setting, and characters from the dataset into the specific data format required by IBSEN.
To achieve this transformation, we employed GPT-4o mini to automatically process the conversion. This step ensures that the data is correctly formatted and fully compatible with IBSEN’s input requirements, ultimately allowing us to generate the necessary script files and other related outputs.
The final output of this method consists of log files containing dialogues between different characters. Since these logs primarily capture character interactions in a structured manner, we leveraged a large language model (LLM) to further refine them into coherent and fluent stories. This post-processing step enhances readability and ensures that the generated stories flow naturally, making them suitable for evaluation and further analysis.
@@
In the multi-agent virtual sandbox simulation of this method, several hyperparameters are configurable to tailor the environment to the needs of the generation process. The simulation is initialized at the timestamp “2024-09-01 12:00 AM”, with each simulation step representing a time interval of one hour. We utilize sqlite3 as the underlying database to store and manage the simulation data.
@@ Appendix I Case Study
‘‘What is this?’’ she murmured, her pulse quickening. The ink was barely legible, but the words ‘time fracture’ and ‘experiment 47’ stood out starkly. [Psychological] She traced the lines with her finger, a mix of excitement and dread bubbling within her. Could this be the key to understanding what had happened?
Just then, Dr. Harold Reed walked in, his presence cutting through the silence like a beacon. ‘‘Ah, Claire. Diving into the past again, are we?’’ His voice carried a warmth that momentarily eased her growing anxiety. ...[246 words]...
@@
As she entered the room, she called out, ‘‘Everyone! I found something! We need to talk about ‘experiment 47’ and what it could mean for us!’’ The world outside was silent,[Psychological] but inside her heart, a fire was beginning to blaze. [Environment] As the team gathered around the long, battered table in the dimly lit makeshift lab, [Psychological] Claire felt a wave of anticipation ripple through her. The notebook lay open before them, its pages a chaotic testament to the desperation that had led to the time freeze. ...[67 words]...
He paused, looking each member of the team in the eye as if urging them to unearth their vulnerabilities. [Environment] The faint hum of the generator in the corner provided an eerie backdrop to the silence that followed, amplifying the gravity of Dr. Reed’s words. ‘‘I propose we share our experiences over the past weeks. Our personal insights might just lead us to the connections we need, both in understanding this phenomenon and in healing ourselves.’’
[Conflict] Maya, sitting cross-legged on her chair, looked up from her sketchbook. ‘‘But Dr. Reed, can we really afford to get sidetracked by our feelings? We need to stay focused on the science!’’ ...[33 words]...
As Maya processed this, Tommy, who had been fidgeting with a piece of scrap metal, finally spoke up. [Conflict] ‘‘I... I can share something. The freeze didn’t just stop time for me; it stopped everything. I used to love scavenging through the old tech, finding treasures in the ruins. But now... it’s like everything I find is just a relic of a past I can’t access. It’s frustrating. I feel like I’m stuck in a still frame.’’ ...[492 words]...
@@
[Psychological] Claire felt a warmth spread through her chest as she watched her teammates collaborate, their energies intertwining like threads in a fabric. ‘‘This is what we need, isn’t it? To remember that even in the face of this freezing silence, we’re still here, still human.’’
@@
[Conflict] Chris, immersed in his laptop, suddenly slammed the lid shut, breaking the silence. ‘‘We can’t keep talking about feelings, Claire! We need data, not emotional fluff! This is about survival, not therapy!’’ His voice rose, echoing off the cold concrete walls, causing heads to turn.
[Psychological] Claire’s heart raced. [Conflict] ‘‘How can you say that? Emotions are part of our experience! We can’t just shove them aside as if they don’t matter! If we don’t understand the human element, how do we hope to solve the anomaly?’’ ...[40 words]...
[Conflict] Chris scoffed, shaking his head. ‘‘You’re both missing the bigger picture! The anomaly is a scientific problem, and we need to treat it as such. Statistics and models, that’s what we need! Not a group therapy session!’’
[Environment] A heavy silence followed his outburst, filled only by the distant hum of the city’s systems, eerily frozen in time. [Psychological] Claire felt a weight settle in her chest, her frustration mixing with the isolation she had been struggling with since the freeze began. She took a deep breath, grounding herself. [Conflict] “I’m not asking for therapy, Chris. I’m asking for understanding. All of us are feeling the strain of this anomaly. If we don’t address the emotional fallout, we risk losing more than just our city—we risk losing ourselves.” ...[39 words]...
[Conflict] Chris threw his hands up, cutting him off. ‘‘That’s ridiculous, Tommy! This isn’t a support group! We’re not here to share our feelings; we’re here to fix a problem!’’ His anger was palpable, and the room tensed at the ferocity of his words. ...[54 words]...
Chris clenched his jaw, visibly wrestling with his emotions as he paced the room, his agitation bouncing off the walls. [Psychological] The others watched, feeling the weight of the moment. Finally, he turned to Claire, his tone softening slightly. “I just... I can’t afford to lose focus. We’re running out of time.”
[Psychological] Claire’s heart softened at his vulnerability. “None of us can afford to lose focus, Chris. But we’re not in this alone, and we shouldn’t have to carry it all on our shoulders. What if we tried combining both approaches? We could analyze the data, but also create a space for sharing. We could document our emotional states alongside our findings. It might help us see patterns we wouldn’t have considered otherwise.” ...[81 words]...
@@
[Environment] As the sun dipped below the horizon, casting a soft golden glow over Frosthaven, Sophia stood in the community center, surveying the room with a mixture of excitement and apprehension. The large, open space was adorned with strings of fairy lights, their warmth contrasting with the cold, sterile environment of their scientific workspace. Canvas draped across tables awaited contributions, and Maya’s masterpiece, ‘Isolation’s Embrace,’ was prominently displayed at the front, its swirling colors capturing the emotional struggles faced by the community. ...[305 words]...
[Conflict] When it was Chris’s turn to speak, he reluctantly approached the makeshift stage, a shadow of frustration crossing his features. ‘‘I appreciate the intent behind this gathering, but we have to remember that feelings don’t solve the problem. We need actionable data—statistics, research, something concrete to work with!’’ ...[405 words]...
[Conflict] Chris sighed, his shoulders slumping. ‘‘I just don’t want to lose focus. If we get too caught up in emotions, we might miss our chance to save the city.’’ ...[35 words]...
@@
[Conflict] ‘‘What have you done?’’ Claire’s voice trembled slightly, a mixture of fear and frustration. She stepped closer, her mind racing with the implications of the power surge.
[Conflict] Chris, his hands still hovering over the keyboard, shot her a defiant glance. ‘‘I was just trying to access the central grid! If we can understand the frozen energy signatures, maybe we can find a way to reverse this freeze. We can save Frosthaven!’’ ...[49 words]...
[Environment] The monitors flickered, reflecting a kaleidoscope of error messages and warnings, each one a testament to the risk they were all contemplating. [Conflict] Tommy leaned over Chris’s shoulder, his eyes darting between the screen and the faces of his friends. ‘‘We should think this through. What if we lose everything?’’ He was usually the one advocating for the thrill of experimentation, but this was different. The stakes felt heavier now.
Sophia, who had been quietly observing, stepped forward. Her voice was steady, an anchor amidst the storm. [Conflict] ‘‘We need to decide if this is the right path. We’re already facing the consequences of our actions—what’s one more gamble on top of all this?’’ Her gaze swept across the group, searching for consensus. ‘‘Maybe there’s another way. Let’s take a moment to regroup and consider all our options.’’
[Psychological] Claire nodded, her heart pounding as she recalled their earlier discussions about balancing emotional clarity with scientific inquiry. ‘‘I agree with Sophia. This isn’t just about reversing the anomaly. It’s about us too. We need to find a solution that doesn’t put any of you at risk.’’ [Conflict] Chris’s face hardened at their resistance, yet a flicker of uncertainty crossed his features. ‘‘But we don’t have time to sit around! Every moment we waste, more lives are impacted. People are trapped in this stasis!’’ His voice rose, desperation lacing his words, revealing the turmoil beneath his bravado. ...[50 words]...
Claire watched Chris as his expression shifted, the wall of bravado beginning to crack. The tension in the room thickened, silence stretching out like an elastic band ready to snap. [Psychological] Finally, Chris slumped back in his chair, his anger dissipating into a weary resignation. ‘‘I just thought... maybe this was the way to prove I’m not just a tech geek who hides behind a screen. I wanted to be part of something bigger.’’ ...[30 words]...
The group sat in contemplative silence, the weight of their shared isolation settling among them. [Psychological] Claire inhaled deeply, her resolve strengthening. ‘‘Let’s take a step back. We can analyze the data from your hack without triggering further consequences. We can create a simulation. Test it. Navigate the risks together.’’
@@
[Environment] As she finished reading, silence enveloped the room, but it wasn’t the same tense silence as before. Instead, it felt warm, inviting a space for reflection. Chris shifted in his seat, his face softening as he took in her words. ‘‘I never thought of it that way, Maya. I’ve been so focused on the data, on proving myself, that I forgot we’re all human here, feeling the same fears. It’s easy to hide behind numbers and screens, but I miss feeling connected.’’ ...[54 words]...
[Psychological] Claire felt a wave of relief wash over her as she realized the significance of Maya’s honesty. ‘‘This is what I’ve been trying to express. Our emotional journeys are just as important as our scientific ones. If we want to move forward, we need to create a safe space where we can be open with each other. We’re not just colleagues; we’re a team that needs to lean on one another.’’ ...[292 words]...
@@
[Environment] When they finally stepped away from the table, the energy in the room felt charged, like the calm before a storm of creativity and collaboration. Claire turned to Dr. Reed, her eyes glinting with determination. ‘‘Let’s bring this to the team. I believe together, we can thaw the emotional freeze and reignite the spirit of Frosthaven.’’

### v2

@@ Introduction
The advent of large language models (LLMs) brings significant advancements to various fields, including multi-agent simulations (Li et al. 2024; Qian et al. 2024; Chen et al. 2024). These simulations offer a powerful tool for modeling complex interactions between virtual agents, providing a dynamic and context-rich environment for story generation. When humans write stories, they typically have an overarching mental picture of the story world. In our approach, multi-agent simulations are used to create this overarching scene, where agents (i.e., the characters in the story) interact based on predefined behaviors, triggering emergent events that form the foundation of the story. This allows for the automatic generation of diverse and engaging scenarios, which are crucial for building long-form stories.
@@
Traditional story generation often follows a top-down method (Zhou et al. 2023; Wang et al. 2023; Fan et al. 2019; Goldfarb-Tarrant et al. 2020; Yang et al. 2022), where a story’s structure is outlined first and then expanded. Although this framework provides a guide, it can limit natural character development and plot progression, resulting in stories that feel forced or predictable. In contrast, our hybrid bottom-up approach starts with agent-driven simulations that generate emergent events. As agents interact, these events unfold naturally, enabling more organic character growth and plot evolution, ultimately producing stories with greater depth, coherence, and spontaneity.
@@
LLM-based multi-agent simulations have gained attention for their advanced language processing and decision-making, enabling nuanced agent interactions (Jinxin et al. 2023; Liu et al. 2023; Feng et al. 2023). These systems have been explored in various domains (Wang et al. 2024a; Williams et al. 2023), particularly social simulation, which is most relevant to our research. Social simulation with LLM-based agents provides an effective way to model complex societal dynamics that are otherwise difficult or costly to study (Li et al. 2023b, a). For example, S3 (Gao et al. 2023) investigates the spread of information, emotions, and attitudes in social networks, while Generative Agents (Park et al. 2023) and AgentSims (Lin et al. 2023) model daily human interactions in virtual towns. Social Simulacra (Park et al. 2022) focuses on simulating online community regulation, and SocialAI School (Kovač et al. 2023) uses LLMs to simulate child development. These studies highlight the growing potential of LLM-based agents to offer valuable insights into societal behavior, community regulation, and social development, which provides a foundation for hybrid bottom-up long-form story generation.
@@
Story generation has advanced with the rise of LLMs, which produce fluent stories but often struggle with coherence and consistency (Huot et al. 2024). To address these challenges, frameworks such as outlining followed by expansion into full stories have been proposed (Zhou et al. 2023; Wang et al. 2023). However, long-form story generation remains challenging. Recent studies have explored multi-agent systems using LLMs (Wang et al. 2024b; Nasir et al. 2024), such as Agents’ Room (Huot et al. 2024), which focuses on generating stories of 1,000-2,000 words using Planning and Writing Agents coordinated by an Orchestrator, and IBSEN (Han et al. 2024), which employs Director-Actor agent collaboration for scriptwriting. While top-down approaches, such as outlining, provide structure, they limit natural storylines and character interactions. In contrast, our hybrid bottom-up approach generates dynamic stories from simulations, producing richer and more coherent stories of over 10,000 words, with clear advantages over traditional methods.
@@
The environment plays a critical role in the sandbox. In Generative Agents, the environment is modeled using a tile-based system, where tiles represent the basic environmental units commonly found in RPG games. However, this approach has inherent limitations, particularly the need for precise coordinates, which can make modeling more rigid and less flexible. To overcome these limitations, we adopt a more general tree-like structure for environment modeling, which does not rely on specific coordinates but instead uses relative distances between objects and areas. For a detailed introduction to environment modeling, please refer to the supplementary material.
@@
Following the setup described in DOC (Yang et al. 2023), we use premises, settings, and characters as inputs. However, the DOC dataset contains relatively homogeneous story types, so we construct a new dataset consisting of 20 story-related settings without reference answers. Our dataset features a broader variety of story types, including genres such as science fiction. For detailed information about our dataset, please refer to the supplementary material.
@@
We conduct structured pairwise comparisons across six criteria: Plot, Creativity, Character Development, Language Use, Conflict Quality, and Overall. These dimensions cover key aspects of long-form storytelling, including narrative coherence, originality, emotional depth, and stylistic quality. Annotators with backgrounds in literature, storytelling, and related fields compare story pairs from the same setting and provide ratings for each dimension as well as an overall preference. To reduce fatigue and bias, each annotator evaluates a limited number of pairs per session. We apply a Latin Square design to balance the assignment of story settings and randomize story order. For each setting, we collect $N\times(N-1)/2$ comparisons, where $N$ is the number of methods.
@@
For full metric definitions, evaluation procedures, and interface details, please refer to the supplementary material.
@@
We select several representative and high-performing methods as our baselines. We categorize the baseline methods into three types: (1) Vanilla LLMs: methods that generate long-form stories directly using large language models, such as GPT-4o (Hurst et al. 2024) and DeepSeek-V3 (Liu et al. 2024); (2) Structured Frameworks: methods that utilize structured frameworks for long-form story generation based on LLMs, such as Re3 (Yang et al. 2022) and DOC-V2 (Yang et al. 2023); and (3) Multi-Agent Simulations: methods that generate through multi-agent simulations, such as IBSEN (Han et al. 2024).
@@ Conclusion
In this paper, we introduced StoryBox, a novel approach for long-form story generation using multi-agent simulations. Our experiments show that StoryBox outperforms existing methods on many metrics. Despite challenges in evaluating story quality, both automatic and human evaluations confirm its effectiveness in generating engaging and coherent stories.
Acknowledgements
The authors are supported by the National Key Research and Development Program of China (Grant No. 2020YFA0712500). The authors would also like to thank National Supercomputer Center in Guangzhou for providing high performance computational resources.
@@ Appendix C Metric Details
  • Character Behavior Consistency: A specialized metric for multi-agent simulation-based story generation, this measures whether characters’ actions align with their predefined persona settings (Persona Scratch Information). Characters should act consistently with their attributes, goals, and motivations, ensuring believability and coherence in the story.
@@ Appendix D Human Evaluation Details
To facilitate consistent and efficient story evaluation, we develop a custom web-based platform tailored to the needs of our human evaluation task. An overview of the platform interface is shown in Figure 8. To accommodate a diverse group of annotators, the platform supports both Chinese and English language modes. Access to the system is restricted to verified annotators, who must log in before participating in the evaluation process to ensure traceability and avoid duplicate submissions.
@@ Appendix F Sandbox Initialization
The setup of sandbox characters is derived from the understanding of the relevant story settings. It primarily involves establishing the “Persona Scratch Information”, as outlined in Figure 2, which includes details such as names, ages, and other relevant information. Based on the description in “Story 1”, we utilize GPT-4o mini to generate the corresponding character setups, as shown in Table LABEL:tab:story_1_character_setup. The same approach is applied to other stories.
@@
The sandbox environment setup is similarly based on the understanding of the relevant story settings. The goal is to construct an environment with hierarchical relationships, as depicted in Figure 7. In this case, drawing from the details in “Story 1”, we use GPT-4o mini to generate the corresponding environment setup in the form of a YAML file, as shown in Table LABEL:tab:story_1_environment_setup. This method is also applied to other stories in the dataset.
@@ Appendix H More Implementation Details
During the reproduction of DOC-V2, we observe that the prompts used in this method often result in parsing failures. Additionally, since this method does not specify the type of story, we set the story type of this method to “narrative” during evaluation. To ensure fairness, we replace all OpenAI engines used in this method with GPT-4o mini.
@@
During the reproduction of IBSEN, we observe that this method primarily utilizes a multi-agent virtual sandbox for generating theatrical scripts. As a result, it is necessary to first convert the premise, setting, and characters from the dataset into the specific data format required by IBSEN.
To achieve this transformation, we employ GPT-4o mini to automatically process the conversion. This step ensures that the data is correctly formatted and fully compatible with IBSEN’s input requirements, ultimately allowing us to generate the necessary script files and other related outputs.
The final output of this method consists of log files containing dialogues between different characters. Since these logs primarily capture character interactions in a structured manner, we leverage a large language model (LLM) to further refine them into coherent and fluent stories. This post-processing step enhances readability and ensures that the generated stories flow naturally, making them suitable for evaluation and further analysis.
@@
In the multi-agent virtual sandbox simulation of this method, several hyperparameters are configurable to tailor the environment to the needs of the generation process. The simulation is initialized at the timestamp “2024-09-01 12:00”, with each simulation step representing a time interval of one hour. We utilize sqlite3 as the underlying database to store and manage the simulation data.
@@ Appendix I Case Study
"What is this?" she murmured, her pulse quickening. The ink was barely legible, but the words "time fracture" and "experiment 47" stood out starkly. [Psychological] She traced the lines with her finger, a mix of excitement and dread bubbling within her. Could this be the key to understanding what had happened?
Just then, Dr. Harold Reed walked in, his presence cutting through the silence like a beacon. "Ah, Claire. Diving into the past again, are we?" His voice carried a warmth that momentarily eased her growing anxiety. ...[246 words]...
@@
As she entered the room, she called out, "Everyone! I found something! We need to talk about ’experiment 47’ and what it could mean for us!" The world outside was silent,[Psychological] but inside her heart, a fire was beginning to blaze. [Environment] As the team gathered around the long, battered table in the dimly lit makeshift lab, [Psychological] Claire felt a wave of anticipation ripple through her. The notebook lay open before them, its pages a chaotic testament to the desperation that had led to the time freeze. ...[67 words]...
He paused, looking each member of the team in the eye as if urging them to unearth their vulnerabilities. [Environment] The faint hum of the generator in the corner provided an eerie backdrop to the silence that followed, amplifying the gravity of Dr. Reed’s words. "I propose we share our experiences over the past weeks. Our personal insights might just lead us to the connections we need, both in understanding this phenomenon and in healing ourselves."
[Conflict] Maya, sitting cross-legged on her chair, looked up from her sketchbook. "But Dr. Reed, can we really afford to get sidetracked by our feelings? We need to stay focused on the science!" ...[33 words]...
As Maya processed this, Tommy, who had been fidgeting with a piece of scrap metal, finally spoke up. [Conflict] "I... I can share something. The freeze didn’t just stop time for me; it stopped everything. I used to love scavenging through the old tech, finding treasures in the ruins. But now... it’s like everything I find is just a relic of a past I can’t access. It’s frustrating. I feel like I’m stuck in a still frame." ...[492 words]...
@@
[Psychological] Claire felt a warmth spread through her chest as she watched her teammates collaborate, their energies intertwining like threads in a fabric. "This is what we need, isn’t it? To remember that even in the face of this freezing silence, we’re still here, still human."
@@
[Conflict] Chris, immersed in his laptop, suddenly slammed the lid shut, breaking the silence. "We can’t keep talking about feelings, Claire! We need data, not emotional fluff! This is about survival, not therapy!" His voice rose, echoing off the cold concrete walls, causing heads to turn.
[Psychological] Claire’s heart raced. [Conflict] "How can you say that? Emotions are part of our experience! We can’t just shove them aside as if they don’t matter! If we don’t understand the human element, how do we hope to solve the anomaly?" ...[40 words]...
[Conflict] Chris scoffed, shaking his head. "You’re both missing the bigger picture! The anomaly is a scientific problem, and we need to treat it as such. Statistics and models, that’s what we need! Not a group therapy session!"
[Environment] A heavy silence followed his outburst, filled only by the distant hum of the city’s systems, eerily frozen in time. [Psychological] Claire felt a weight settle in her chest, her frustration mixing with the isolation she had been struggling with since the freeze began. She took a deep breath, grounding herself. [Conflict] "I’m not asking for therapy, Chris. I’m asking for understanding. All of us are feeling the strain of this anomaly. If we don’t address the emotional fallout, we risk losing more than just our city—we risk losing ourselves." ...[39 words]...
[Conflict] Chris threw his hands up, cutting him off. "That’s ridiculous, Tommy! This isn’t a support group! We’re not here to share our feelings; we’re here to fix a problem!" His anger was palpable, and the room tensed at the ferocity of his words. ...[54 words]...
Chris clenched his jaw, visibly wrestling with his emotions as he paced the room, his agitation bouncing off the walls. [Psychological] The others watched, feeling the weight of the moment. Finally, he turned to Claire, his tone softening slightly. "I just... I can’t afford to lose focus. We’re running out of time."
[Psychological] Claire’s heart softened at his vulnerability. "None of us can afford to lose focus, Chris. But we’re not in this alone, and we shouldn’t have to carry it all on our shoulders. What if we tried combining both approaches? We could analyze the data, but also create a space for sharing. We could document our emotional states alongside our findings. It might help us see patterns we wouldn’t have considered otherwise." ...[81 words]...
@@
[Environment] As the sun dipped below the horizon, casting a soft golden glow over Frosthaven, Sophia stood in the community center, surveying the room with a mixture of excitement and apprehension. The large, open space was adorned with strings of fairy lights, their warmth contrasting with the cold, sterile environment of their scientific workspace. Canvas draped across tables awaited contributions, and Maya’s masterpiece, "Isolation’s Embrace," was prominently displayed at the front, its swirling colors capturing the emotional struggles faced by the community. ...[305 words]...
[Conflict] When it was Chris’s turn to speak, he reluctantly approached the makeshift stage, a shadow of frustration crossing his features. "I appreciate the intent behind this gathering, but we have to remember that feelings don’t solve the problem. We need actionable data—statistics, research, something concrete to work with!" ...[405 words]...
[Conflict] Chris sighed, his shoulders slumping. "I just don’t want to lose focus. If we get too caught up in emotions, we might miss our chance to save the city." ...[35 words]...
@@
[Conflict] "What have you done?" Claire’s voice trembled slightly, a mixture of fear and frustration. She stepped closer, her mind racing with the implications of the power surge.
[Conflict] Chris, his hands still hovering over the keyboard, shot her a defiant glance. "I was just trying to access the central grid! If we can understand the frozen energy signatures, maybe we can find a way to reverse this freeze. We can save Frosthaven!" ...[49 words]...
[Environment] The monitors flickered, reflecting a kaleidoscope of error messages and warnings, each one a testament to the risk they were all contemplating. [Conflict] Tommy leaned over Chris’s shoulder, his eyes darting between the screen and the faces of his friends. "We should think this through. What if we lose everything?" He was usually the one advocating for the thrill of experimentation, but this was different. The stakes felt heavier now.
Sophia, who had been quietly observing, stepped forward. Her voice was steady, an anchor amidst the storm. [Conflict] "We need to decide if this is the right path. We’re already facing the consequences of our actions—what’s one more gamble on top of all this?" Her gaze swept across the group, searching for consensus. "Maybe there’s another way. Let’s take a moment to regroup and consider all our options."
[Psychological] Claire nodded, her heart pounding as she recalled their earlier discussions about balancing emotional clarity with scientific inquiry. "I agree with Sophia. This isn’t just about reversing the anomaly. It’s about us too. We need to find a solution that doesn’t put any of you at risk." [Conflict] Chris’s face hardened at their resistance, yet a flicker of uncertainty crossed his features. "But we don’t have time to sit around! Every moment we waste, more lives are impacted. People are trapped in this stasis!" His voice rose, desperation lacing his words, revealing the turmoil beneath his bravado. ...[50 words]...
Claire watched Chris as his expression shifted, the wall of bravado beginning to crack. The tension in the room thickened, silence stretching out like an elastic band ready to snap. [Psychological] Finally, Chris slumped back in his chair, his anger dissipating into a weary resignation. "I just thought... maybe this was the way to prove I’m not just a tech geek who hides behind a screen. I wanted to be part of something bigger." ...[30 words]...
The group sat in contemplative silence, the weight of their shared isolation settling among them. [Psychological] Claire inhaled deeply, her resolve strengthening. "Let’s take a step back. We can analyze the data from your hack without triggering further consequences. We can create a simulation. Test it. Navigate the risks together."
@@
[Environment] As she finished reading, silence enveloped the room, but it wasn’t the same tense silence as before. Instead, it felt warm, inviting a space for reflection. Chris shifted in his seat, his face softening as he took in her words. "I never thought of it that way, Maya. I’ve been so focused on the data, on proving myself, that I forgot we’re all human here, feeling the same fears. It’s easy to hide behind numbers and screens, but I miss feeling connected." ...[54 words]...
[Psychological] Claire felt a wave of relief wash over her as she realized the significance of Maya’s honesty. "This is what I’ve been trying to express. Our emotional journeys are just as important as our scientific ones. If we want to move forward, we need to create a safe space where we can be open with each other. We’re not just colleagues; we’re a team that needs to lean on one another." ...[292 words]...
@@
[Environment] When they finally stepped away from the table, the energy in the room felt charged, like the calm before a storm of creativity and collaboration. Claire turned to Dr. Reed, her eyes glinting with determination. "Let’s bring this to the team. I believe together, we can thaw the emotional freeze and reignite the spirit of Frosthaven."